3. In a random sample of 193 engineers, 15.2% of them were women. a) Construct an 85% confidence interval for the populationPlease Show Work. Thanks.
Answer #1

Solution :

Given that,

a) Point estimate = sample proportion = \hat p = 0.152

1 - \hat p = 1 - 0.152 = 0.848

Z\alpha/2 = Z0.075 = 1.44

Margin of error = E = Z\alpha / 2 * ((\hat p * (1 - \hat p ))\sqrt / n)

= 1.44 (\sqrt((0.152 * 0.848) / 193)

= 0.037

A 85% confidence interval for population proportion p is ,

\hat p ± E   

= 0.152  ± 0.037

= ( 0.115, 0.189 )

= ( 11.5%, 18.9% )

We are 85% confident that true proportion of them were women between 11.5% and 18.9%

b) Margin of error = E = Z\alpha / 2 * ((\hat p * (1 - \hat p ))\sqrt / n)

= 1.44 (\sqrt((0.152 * 0.848) / 193)

= 0.037

= 3.7%